OpenStudy (anonymous):

Susan's math teacher assigned x homework problems on Monday. On Tuesday, she assigned 15 more problems. Over the two days, a total of 25 homework problems were assigned. Which equation could be used to find x, the number of problems assigned on Monday?

OpenStudy (anonymous):

A. 15x = 25 x + 15 = 25 25 + x = 15 x - 15 = 25